Assuming that a force of 100 N is pulling an object across a surface (with friction) at a speed of 20 m/s to the right, while a friction force of 100 N magnitude is causing a drag force to the left, what is the speed of the object 10 seconds later?
1
Expert's answer
2012-10-30T11:08:38-0400
The net force acting on an object is

F(pulling) - F(friction) = 100[N] - 100[N] = 0[N].

According to the second Newton's law, the acceleration of an object has zero value:

a = F/m = 0/m = 0.

Therefore, speed of an object remains constant. So, 10 seconds later the speed of an object is 20 m/s.